OpenStudy (anonymous):

Which pair of ratios forms a proportion? 3 over 16. and 6 over 32. 32 over 11. and 16 over 22. 10 over 22. and 20 over 11. 14 over 4. and 28 over 2.

OpenStudy (anonymous):

good job omistone

OpenStudy (anonymous):

well for example 1/2 and 4/8 are a proportion :)

OpenStudy (anonymous):

the best way is to evaluate each one. Lets start with A 3/16=6/32 both are .1875 so it is proportionate .1875=.1875 this is your answer =================== B is 32 / 11= 16 / 22 2.909=.72727 when you evaluate aka type into a calculator you get Which is not true not proportionate

OpenStudy (anonymous):

B isn't .

OpenStudy (anonymous):

A :P i told you silly

OpenStudy (anonymous):

Ohh , loll .

OpenStudy (anonymous):

Thanks . I understand it now .

OpenStudy (anonymous):

LIKe i said 1/2 and 4/8 are a proportion. Plug each one to a calculator .5 and .5 .5=.5 therefore a proportion. ================= A 3/16=6/32 both are .1875 so it is proportionate .1875=.1875 this is your answer --------------------------- B 32 / 11= 16 / 22 2.909=.72727 not true not a proportion.